OpenStudy (anonymous):

help indicate which type of conic the following represents (x-3)^2/9+(y-2)^2/4=1

OpenStudy (anonymous):

ellipse

OpenStudy (harsimran_hs4):

x^2/a + y^2/b = 1 is general eq of ellipse so x-3 and y-2 just represents shifting of the center of ellipse

OpenStudy (anonymous):

how would we know that? i mean i have other prob;ems like that but u kno hw u said elipse how can i figure it out

OpenStudy (anonymous):

oh really that is? i didmt know that

OpenStudy (anonymous):

(x=1)^2/25+(y-2)^2/25=1 would also be an elipse then right?

OpenStudy (harsimran_hs4):

@nisa http://en.wikipedia.org/wiki/Ellipse read about elements of the ellipse

OpenStudy (harsimran_hs4):

yes if it`s x-1 instead of x = 1

OpenStudy (anonymous):

alright i will read that thank you so the second one that i said would not be an elipse unless it was x-1 which its not

OpenStudy (harsimran_hs4):

look to get hold of this concept just learn 2 things 1. definition of the ellipse, this will help you right the general equation of ellipse 2. shifting of the axes/ origin so when ever you notice things like x-a instead of x where a is constant it just means that the curve shifted the position but the shape does not change at all

OpenStudy (anonymous):

and x is the shape

OpenStudy (harsimran_hs4):

*write so you getting some idea about it ?

OpenStudy (anonymous):

and anything plus or minus it is just a curve or a change but not in the shape

OpenStudy (harsimran_hs4):

any + or - constant is just shifting of curve with no change in shape

OpenStudy (anonymous):

alright got it thank you :)

OpenStudy (harsimran_hs4):

great!!

OpenStudy (anonymous):

i really appreciate it :)